OpenStudy (anonymous):

n a class of 50 students, the number of women is 2 more than five times the number of men. How many women are in the class?

OpenStudy (anonymous):

m + w = 50 and w = 5m + 2 These are 2 simultaneous equation in two variables. Can be solved now by either substitution or elimination.

OpenStudy (anonymous):

m + w = 50 w = 5m +2 m=8 w=42

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Thanks It worked out